OpenStudy (anonymous):

PLZ HELP NOW!!!! Xavier's dining room table was unstable because one leg was shorter than the others. He placed a 3/8 inch thick piece of wood under the short leg, and it became even with the other legs, which are 19 1/4 inches long. How many inches long was the short leg before Xavier added the wood? 19 5/8 inches 18 3/4 inches 19/ 7/8 inches 18 7/8 inches

OpenStudy (anonymous):

@angel12310

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

hi

OpenStudy (anonymous):

can u plz help me with this question

OpenStudy (anonymous):

sure i'll see

OpenStudy (anonymous):

18 7/8

OpenStudy (anonymous):

3 legs are length 19 1/4 but one leg is shorter by 3/8 its length is then 19 1/4 - 3/8 = 18 7/8 inches

OpenStudy (anonymous):

anymore

OpenStudy (anonymous):

YUP WAIT THANKS

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

THANKS FOR THE MEDAL, HOW CAN I GIVE U A MEDAL?

OpenStudy (anonymous):

click best response near my name

OpenStudy (anonymous):

anymore questions though just type them

OpenStudy (anonymous):

I DID

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

but have you gotten one from me?

OpenStudy (anonymous):

what do you mean if medal then yes

OpenStudy (anonymous):

questions

OpenStudy (anonymous):

i'm gonna go to another question ok while you think

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

think quick or i'll have to leave

OpenStudy (anonymous):

no wait

OpenStudy (anonymous):

nevermind i have no more bye thk

OpenStudy (anonymous):

bye thank you

OpenStudy (anonymous):

i was joking i'm not leaving take your time to think

OpenStudy (anonymous):

talk to you tomorrow bye

OpenStudy (anonymous):

omg really lol

OpenStudy (anonymous):

LOL

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok wait i'm looking for the ques

OpenStudy (anonymous):

Solve 1/6h - 2/3 = -1/4 for h h= 2 1/2 h= -11/72 h= -5 1/2 h= 5/72

OpenStudy (anonymous):

@angel12310?

OpenStudy (anonymous):

help

OpenStudy (anonymous):

help

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

i can't do this one it is hard

OpenStudy (anonymous):

Solve 1/6h - 2/3 = -1/4 for h h= 2 1/2 h= -11/72 h= -5 1/2 h= 5/72

OpenStudy (anonymous):

Solve 1/6h - 2/3 = -1/4 for h h= 2 1/2 h= -11/72 h= -5 1/2 h= 5/72

OpenStudy (anonymous):

i said i can't do it

OpenStudy (anonymous):

sorry

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

ok